OpenStudy (anonymous):

Help please..For the function y=1^x, make a table with integer values of x from 0 to 4. Then graph the function (you do not need to submit the graph). Does the graph of show exponential growth, exponential decay, or neither? Explain your thinking.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

y = 1^x x = 0, y = 1^0 = 1 (anything raised to 0 is 1) x = 1, y = 1^1 = 1 x = 2, y = 1^2 = 1 x = 3, y = 1^3 = 1 x = 4, y = 1^4 = 1 The y value remains the same for all x values. Does the graph of show exponential growth, exponential decay, or neither? Answer: Neither Explain your thinking. Answer: The y value remains the same for all x values. So it is neither growing nor decaying.

OpenStudy (anonymous):

thank you so much

OpenStudy (ranga):

u r welcome.

OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):



OpenStudy (ranga):

what is the definition of tan for angle C?

OpenStudy (anonymous):

30

OpenStudy (ranga):

angle C is 30 degrees. Are you familiar with trigonometric function such as sin, cos, tan of an angle?

OpenStudy (anonymous):

not really, it confuses me